How long does it take to double an investment of $20680 if the investment pays only simple interest at the rate of 10% per year? Enter your answer as: 1.23 Round your answer to 2 decimal places (Canvas will add 2 extra decimal places, but ignore those
Answer #1

Simple interest formula is

SI=P*R*t divided by 100

and Amount = principle (P) + simple interest (SI)

when amount is doubled.i.e. 2P then

SI= 2P-P = P

and P = P*10*t divided by 100

implies t= 10 years